# Fast Jet Tagging with MLP-Mixers on FPGAs

### v1

@@ 1 Introduction
The CERN Large Hadron Collider (LHC) [1] generates vast quantities of data, producing hundreds of terabytes per second from proton-proton collisions occurring every 25 ns. The first reduction of this enormous data throughput is managed by the hardware-based level-1 trigger system (L1T), which filters events in real-time with stringent latency requirements of a few microseconds [2, 3]. This system, utilizing approximately 1,000 Field-Programmable Gate Arrays (FPGAs), ensures critical events are preserved for offline analysis while drastically reducing the data rate by two orders of magnitude, thereby managing the downstream bandwidth effectively. In this context, the algorithms’ accuracy is crucial while the computational complexity of each algorithm must be kept minimal as the resources are scarce. The forthcoming High-Luminosity LHC (HL-LHC) [4] upgrade will significantly increase the data rate and complexity, posing challenges for real-time data processing. Machine learning techniques are being actively explored to enhance the speed and accuracy of the trigger algorithms [2, 3]; however, addressing strict resource and latency constraints without compromising performance present important challenges.
@@
Jet tagging is a crucial task for physics analyses at the LHC, which enables the identification of rare events that probe fundamental questions about the universe, such as the existence of new particles or the verification of theoretical models. This task involves identifying the origin of jets, which are collimated sprays of particles produced in high-energy collisions, and associating them with specific underlying physics processes. While jet tagging is traditionally performed offline using large machine learning models and high-performance computing resources [5, 6, 7, 8, 9], extending this capability to the trigger system ensures that such valuable events are not discarded prematurely. However, this transition is highly challenging due to the stringent constraints of the trigger system. Offline models, designed for accuracy without hardware constraints, often exceed the memory, processing speed, and on-chip resource availability of FPGAs, necessitating significant architectural adaptations and optimizations to make them viable for real-time deployment.
@@
MLP-Mixer [10] is an architecture consisting solely of Multi-layer Perception (MLP) layers, originally designed for vision tasks. For those applications, it splits the input into patches and processes them as a sequence of vectors. As the inputs for jet-tagging tasks at the trigger level are already sequences of particles constituents, MLP-Mixer offers a plausible structure for online jet tagging tasks. In this work, we explore and demonstrate the application of MLP-Mixers for jet-tagging, as well as their deployment on FPGAs for use in low-level trigger systems. We train the models with High-Granularity Quantization (HGQ) [11] to fit within FPGA constraints, and further optimizes their firmware with Distributed Arithmetic (DA) to minimize the latency and maximizing throughput. The firmware is generated and deployed using the hls4ml [12] framework with the Vitis HLS [13] backend, ensuring compatibility and efficiency on FPGA hardware.
@@
  • We demonstrate the deployment of MLP-Mixer models on FPGAs, achieving low latency and high throughput, meeting the requirements of the low-level trigger system.
  • We highlight that MLP-Mixer models outperform prior architectures in accuracy while using significantly fewer resources, achieving higher throughput, and lower latency.
  • We emphasize the advantage of using a non-permutation-invariant model for jet tagging when leveraging HGQ, which allows for selective feature prioritization via heterogeneous activation bitwidths to minimize resource consumption.
@@ 2.1 Fast Jet tagging on FPGAs
Jet tagging is one of the most critical and well-studied tasks in high-energy physics experiments, with various machine learning algorithms proposed to address it [5, 6, 7, 8, 9, 14]. State-of-the-art models for jet tagging include those based on Graph Neural Networks (GNN), such as Particle Net [5] and PELLICAN [7], or Transformer architectures like Particle Transformer [6] and the Lorentz-Equivariant Geometric Algebra Transformers [8].
While these models demonstrate excellent performance for offline data processing, they are typically too large and computationally intensive for real-time deployment on FPGA-based trigger systems. Research efforts [15, 16, 17, 18, 19, 12] have focused on adapting machine learning models for FPGA deployment in real-time environments, showcasing their potential for ultra-low latency and high throughput inference. Notably, Ref. [16] demonstrated that quantization-aware training (QAT) could enable jet tagging with quantized neural networks on FPGAs, achieving competitive accuracy to the full precision models with significantly reduced resource usage. However, these efforts often relied on toy datasets with high-level jet features which are unavailable in real-time trigger systems, making them impractical for actual deployment.
For real-time jet tagging on FPGAs, models like Particle Net and Particle Transformer are quickly ruled out due to their size and complexity. More resource-efficient alternatives include JEDI-net [9], an efficient architecture based on Interaction Networks (IN) [20], and models using the Deep Sets (DS) architecture [21], are being explored. JEDI-net has been successfully implemented on FPGAs [22] with realistic latency and resource consumption, while the DS-based model achieved better latency and resource utilization than IN-based models under comparable hardware-aware constraints [23]. These advancements highlight the feasibility of deploying highly-performant particle-based machine learning models on FPGAs for jet tagging, but further innovations are needed to balance performance, resource efficiency, and latency.
@@
hls4ml [12] is a framework that translates trained neural networks into C++ based high-level synthesis (HLS) projects, which can be further compiled by HLS backends into description languages such as Verilog or VHDL. It supports a wide range of neural network architectures and HLS backends. When the input neural network is properly quantized and configured, hls4ml may generate firmware that is bit-accurate with the original model. This framework has been widely used for deploying neural networks on the hardware triggers at the LHC [16, 17, 15, 24, 25, 26, 27, 2, 3]. Notably, it has been employed for real-time data processing in the L1 trigger system of the CMS experiment, demonstrating its reliability and effectiveness [26, 27]. In this case, hls4ml has facilitated the implementation of autoencoders based anomaly detection models on FPGAs for physics-based applications, achieving remarkable success during LHC Run-3 data taking.
In this work, we leverage hls4ml with the HGQ [11] front-end and the Vitis HLS backend to generate the firmware for the MLP-Mixer models. This combination ensures bit-accuracy between the quantized python model and the generated firmware, while optimizing for resource efficiency on FPGAs.
@@ 2.3 MLP-Mixer
MLP-Mixer [10] is a neural network architecture that exclusively uses Multi-layer Perception (MLP) layers, which is originally designed for vision tasks. For a vision task, the model first splits the image into patches, and treat the flattened vector from each patch as the feature embeddings. The core architecture of the networks involves a series of MLP layers applied along orthogonal directions to mix both spatial- and channel-wise information. Despite its simplicity, this approach has shown competitive performance in computer vision tasks. Although initially developed for image data, we found the MLP-Mixer architecture is well-suited for jet tagging tasks in high-energy physics. Since the input to the jet tagging model is already a sequence of vectors of particle features at L1 triggers, we directly apply the MLP-Mixer architecture on this data without additional modifications. The absence of complex mechanisms, such as self-attention in Transformers involving multiplication of variable-to-variable multiplication operations, simplifies the conversion of MLP-Mixer models with hls4ml and therefore their deployment on FPGAs and its optimization with DA.
@@ 3.1 Dataset
We use the hls4ml jet tagging dataset [28] for training and evaluation. This dataset comprises 620,000 jets in the training set and 260,000 jets in the validation set. The dataset is balanced across 5 classes: Jets originating from gluons (g), light quarks (q), W bosons (W), Z bosons (Z), and top quarks (t). Each jet contains up to 150 particles. Each particle is characterized by 16 kinematic features, as documented in Table 1. The distributions of particle count for each class are shown in Figure I, with the average number of particles per jet varying from 38 to 67, depending on the class.
Table 1: Description of the 16 kinematic features used for each particle in the hls4ml jet tagging dataset.
@@
$p_{z}$ The $x$ component of the momentum of the particle.
@@
$E_{\mathrm{rel}}$ $E/E_{\mathrm{jet}}$, relative energy of the particle.
@@
$\eta_{\mathrm{rot}}$ $\eta$ rotated as described in [29]
@@
$\phi_{\mathrm{rot}}$ $\phi$ rotated as described in [29]
@@
Furthermore, one may apply a typical preselection requirement of a minimum transverse momentum ($p_{T}$) of 2 GeV per particle to simulate the constraints of the trigger system at HL-LHC [30]. With this $p_{T}$ threshold, the average particle count per jet decreases to a range of 31 to 49 as shown in Figure I. If a jet contains fewer than $N$ particles, we pad the missing features with zeros.
@@
At the CMS Experiment [2], we expect the information of up to 128 particles to be available in the trigger system [30]. However, the number of particles required for downstream tasks, varies depending on the application. For instance, while it has been shown that a good performance in b-tagging requires only $\sim 10$ particles [31], we have observed that the jet-tagging task on this dataset benefits from a larger number of particles in the input. Also, event-level tasks, such as missing transverse energy estimation or anomaly detection, may benefit from a larger input to represent the full event. To account for these variations, we train the models with 16, 32, 64, and 128 particles to cover the expected range of inputs.
@@ 3.2 Model
We adopt the MLP-Mixer architecture proposed in [10] as the backbone of our model. However, since the original model is too large for practical FPGA implementation, we designed a downscaled version with only two mixer stages and a dense classification head. Nevertheless, we find that this simplified model is still capable to achieve state-of-art-performance. The exact architecture is depicted in Figure II. Following [22], we use all 16 particle features available in the dataset, and train the models with varying number of particles: 16, 32, 64, and 128.
@@
In contrast to [22], we do not enforce permutation invariance on our models, and all inputs are ordered by $p_{T}$. This design choice is made for several reasons:
@@
Figure II: Architecture of the MLP-Mixer models used in this work. Each model consists of four MLP blocks with a single skip-connection. The input channel size match the number of per-particle input features. DenseBn represents a dense layer followed by a batch normalization layer during training, which are fused into a single layer during inference.
@@ 3.3 Training
All models are implemented and trained with HGQ [11] on TensorFLow 2.13. We use the Adam optimizer with an initial learning rate of $5\cdot 10^{-3}$, along with the CosineDecayRestarts scheduler that resets the learning rate every 100 epochs. The full-precision models are trained for 500 epochs with a batch size of 512 with the categorical cross-entropy loss. The models are trained on a single NVIDIA GTX 1080 GPU with 8GB of memory. The training time for each model ranges between approximately 90 and 120 minutes [^1] Only as a reference. After code optimization, it was found that the training time can be reduced by a factor of 10., depending on the number of particles in the input. Only the models on the Pareto Frontier of the best validation accuracy and resource consumption estimated by Effective Bit Operations (EBOPs) [11] are selected and used to evaluate the performance on the test dataset.
@@ 4.1.1 High Granularity Quantization
We adopt the HGQ [11] method to train, quantize, and prune the MLP-Mixer models. HGQ is an advanced QAT method that optimizes the bitwidths of each individual parameter of the neural network with surrogate gradients on bitwidths. Comparing to traditional QAT methods that optimize the bitwidths at the layer level, such as QKeras [16] or Brevitas [32], HGQ provides a more fine-grained optimization that can achieve significant resource usage reduction for a fully unrolled firmware design while maintaining the accuracy of the original full-precision model. As HGQ allow the bitwidths to reach zero, unstructured pruning is naturally provided. HGQ is a crucial component of our workflow, which ensures that the MLP-Mixer models can be successfully deployed on FPGAs, as synthesis becomes infeasible if the model is not sufficiently quantized.
@@ 4.1.2 Distributed Arithmetic
DA [33, 34, 35, 36, 37] is a method that transforms a linear digital signal processing (DSP) problem with one constant term, such as finite impulse response filters, discrete sine transform, or discrete cosine transform, into a series of operations implemented without the need for dedicated multipliers. This typically involves decomposing constant multiplication operations into a sequence of additions/subtractions and bit-shifts. Optimizations can be applied to the DA implementation to reduce the number of operations and the resource usage. In this work, we adopt a variant of the constant matrix-vector multiplication (CMVM) algorithm proposed in [34] to optimize the CMVM operations in the MLP-Mixer models. An outline of the optimization is as follows:
@@
We use the da4ml [38] library as an extension to the hls4ml library to apply this optimization. By applying the stated DA optimization, we significantly reduce the latency and initiation interval of the models.
@@ 4.2 Training
Following [11], we use a single training for each model to cover the entire Pareto frontier between accuracy and resource consumption. We initialize the models with a small $\beta$ – a scalar parameter that encourages smaller models in a hardware resource-aware approach – and increase its value during training with an exponential scheduler. During training, we save all models that lie on the Pareto frontier between validation accuracy and the resource consumption estimated by the approximated Effective Bit-Operations ($\overline{\mbox{EBOPs}}$[11] provided by the HGQ framework. All other hyperparameters remain identical to those used in the full-precision training. Each model is trained for 5,000 epochs, requiring approximately 22 hours on a single NVIDIA GTX 1080 GPU. However, note that this lengthy training is only required for covering the entire Pareto frontier, and the training time can be significantly reduced if only one performing model is required. For instance, all best-performing quantized models depicted in this work are obtained within 400 epochs.

### v2

@@ 1 Introduction
The CERN Large Hadron Collider (LHC) [lhc1995large] generates vast quantities of data, producing hundreds of terabytes per second from proton-proton collisions occurring every 25 ns. The first reduction of this enormous data throughput is managed by the hardware-based level-1 trigger system (L1T), which filters events in real-time with stringent latency requirements of a few microseconds [cms-tdr-021, atlas-tdr-029]. This system, utilizing approximately 1,000 Field-Programmable Gate Arrays (FPGAs), ensures critical events are preserved for offline analysis while drastically reducing the data rate by two orders of magnitude, thereby managing the downstream bandwidth effectively. In this context, the algorithms’ accuracy is crucial while the computational complexity of each algorithm must be kept minimal as the resources are scarce. For each individual algorithm, the latency requirement could range from a few microseconds down to a few hundred nanoseconds depending on the specific task. A throughput of 40 MHz, the rate at which the L1T receives data, is required for all algorithms. The forthcoming High-Luminosity LHC (HL-LHC) [hl-lhc] upgrade will significantly increase the data rate and complexity, posing challenges for real-time data processing. Machine learning techniques are being actively explored to enhance the speed and accuracy of the trigger algorithms [cms-tdr-021, atlas-tdr-029]; however, addressing strict resource and latency constraints without compromising performance present important challenges.
@@
Jet tagging is a crucial task for physics analyses at the LHC, which enables the identification of rare events that probe fundamental questions about the universe, such as the existence of new particles or the verification of theoretical models. This task involves identifying the origin of jets, which are collimated sprays of particles produced in high-energy collisions, and associating them with specific underlying physics processes. While jet tagging is traditionally performed offline using large machine learning models and high-performance computing resources [pnet, part, pelican, lagtr, jedi-net], extending this capability to the L1T ensures that such valuable events are not discarded prematurely. However, this transition is highly challenging due to the stringent constraints of the trigger system. Offline models, designed for accuracy without hardware constraints, often exceed the memory, processing speed, and on-chip resource availability of FPGAs, necessitating significant architectural adaptations and optimizations to make them viable for real-time deployment.
@@
MLP-Mixer [mlp-mixer] is an architecture consisting solely of Multi-layer Perception (MLP) layers, originally designed for vision tasks. For those applications, it splits the input into patches and processes them as a sequence of vectors. As the inputs for jet-tagging tasks at the trigger level are already sequences of particles constituents, MLP-Mixer offers a plausible structure for online jet tagging tasks. In this work, we explore and demonstrate the application of MLP-Mixers for jet-tagging, as well as their deployment on FPGAs for use in low-level trigger systems. We train the models with High-Granularity Quantization (HGQ) [hgq] to fit within FPGA constraints, and further optimizes their firmware with Distributed Arithmetic (DA) to minimize the latency and maximizing throughput. The firmware is generated and deployed using the hls4ml [hls4ml] framework with the Vitis HLS [vitis] backend, ensuring compatibility and efficiency on FPGA hardware.
@@
  • We demonstrate the deployment of MLP-Mixer models on FPGAs, achieving low latency and high throughput, meeting the requirements of the L1T system at the HL-LHC.
  • We highlight that MLP-Mixer models outperform prior architectures in accuracy while using significantly fewer resources, achieving higher throughput, and lower latency on FPGAs.
  • We emphasize the advantage of using a non-permutation-invariant model for jet tagging when leveraging HGQ, which allows for particle-wise selective feature prioritization via heterogeneous activation bitwidths to minimize resource consumption.
@@ 2.1 Fast Jet tagging on FPGAs
Jet tagging is one of the most critical and well-studied tasks in high-energy physics experiments, with various machine learning algorithms proposed to address it [pnet, part, pelican, lagtr, jedi-net, efn]. State-of-the-art models for jet tagging include those based on Graph Neural Networks (GNN), such as Particle Net [pnet] and PELLICAN [pelican], or Transformer architectures like Particle Transformer [part] and the Lorentz-Equivariant Geometric Algebra Transformers [lagtr].
While these models demonstrate excellent performance for offline data processing, they are typically too large and computationally intensive for real-time deployment on FPGA-based trigger systems. Research efforts [tgc, qkeras, qkeras-xtre-q, autoq, hls4ml, llp, compress-pipe] have focused on adapting machine learning models for FPGA deployment in real-time environments, showcasing their potential for ultra-low latency and high throughput inference. Notably, Ref. [qkeras] demonstrated that quantization-aware training (QAT) could enable jet tagging with quantized neural networks on FPGAs, achieving competitive accuracy to the full precision models with significantly reduced resource usage. However, these efforts often relied on toy datasets with high-level jet features which are unavailable in real-time trigger systems, making them impractical for actual deployment.
For real-time jet tagging on FPGAs, models like Particle Net and Particle Transformer are quickly ruled out due to their size and complexity. More resource-efficient alternatives include JEDI-net [jedi-net], an efficient architecture based on Interaction Networks (IN) [in], and models using the Deep Sets (DS) architecture [ds], are being explored. JEDI-net has been successfully implemented on FPGAs [jedi-fpga] with realistic latency and resource consumption, while the DS-based model achieved better latency and resource utilization than IN-based models under comparable hardware-aware constraints [ds-fpga]. These advancements highlight the feasibility of deploying highly-performant particle-based machine learning models on FPGAs for jet tagging, but further innovations are needed to balance performance, resource efficiency, and latency.
@@
hls4ml [hls4ml] is a framework that translates trained neural networks into C++ based high-level synthesis (HLS) projects, which can be further compiled by HLS backends into description languages such as Verilog or VHDL. It supports a wide range of neural network architectures and HLS backends. When the input neural network is properly quantized and configured, hls4ml may generate firmware that is bit-accurate with the original model. This framework has been widely used for deploying neural networks on the hardware triggers at the LHC [qkeras, qkeras-xtre-q, tgc, ae-l1t, llpnet, axol1tl, cicada, cms-tdr-021, atlas-tdr-029]. Notably, hls4ml has been employed for implementing autoencoder-based anomaly detection triggers [axol1tl, cicada] in the L1T system of the CMS experiment, demonstrating its reliability and effectiveness on FPGAs for physics-based applications and achieving remarkable success during LHC Run-3 data taking.
In this work, we leverage hls4ml with the HGQ [hgq] front-end and the Vitis HLS backend to generate the firmware for the MLP-Mixer models. This combination ensures bit-accuracy between the quantized python model and the generated firmware, while optimizing for resource efficiency on FPGAs.
@@ 2.3 MLP-Mixer
MLP-Mixer [mlp-mixer] is a neural network architecture that exclusively uses Multi-layer Perception (MLP) layers, which is originally designed for vision tasks. For a vision task, the model first splits the image into patches, and treat the flattened vector from each patch as the feature embeddings. The core architecture of the networks involves a series of MLP layers applied along orthogonal directions to mix both spatial- and channel-wise information. Despite its simplicity, this approach has shown competitive performance in computer vision tasks. Since the input to the jet tagging model is already a sequence of vectors of particle features at L1 triggers, we directly apply the MLP-Mixer architecture on this data without additional modifications. The absence of complex mechanisms, such as self-attention in Transformers involving multiplication of variable-to-variable multiplication operations, simplifies the conversion of MLP-Mixer models with hls4ml and therefore their deployment on FPGAs and its optimization with DA.
@@ 3.1 Dataset
We use the hls4ml jet tagging dataset [hls4ml-dataset] for training and evaluation. The dataset contains simulated jets originated from p-p collisions at $\sqrt{s}=13$ TeV. Five classes of jets are included: Jets originating from gluons (g), light quarks (q), W bosons (W), Z bosons (Z), and top quarks (t). All these jets are originated from di-jet events, where both patron or undecayed gauge bosons are generated with $p_{T}\approx 1$ TeV. Together with the underlying event, the particles are clustered with the AK8 algorithm. This dataset comprises 620,000 jets in the training set and 260,000 jets in the validation set, both balanced across the five classes. Each jet contains up to 150 particles, and each particle is characterized by 16 kinematic features, as documented in Table 1. The distributions of particle count for each class are shown in Figure I, with the average number of particles per jet varying from 38 to 67, depending on the class.
Table 1: Description of the 16 kinematic features used for each particle in the hls4ml jet tagging dataset. For the exact definition of each feature, please refer to [hls4ml-dataset].
@@
$p_{z}$ The $z$ component of the momentum of the particle.
@@
$_{\mathrm{rel}} E/E$$E_{\mathrm{jet}}$, relative energy of the particle.
@@
$\eta_{\mathrm{rot}}$ $\eta$ rotated as described in [rot]
@@
$\phi_{\mathrm{rot}}$ $\phi$ rotated as described in [rot]
@@
Furthermore, one may apply a typical preselection requirement of a minimum transverse momentum ($p_{T}$) of 2 GeV per particle to simulate the constraints of the trigger system at HL-LHC [SCRecoL1Jet]. With this $p_{T}$ threshold, the average particle count per jet decreases to a range of 31 to 49 as shown in Figure I. If a jet contains fewer than $N$ particles, we pad the missing features with zeros.
@@
At the CMS Experiment [cms-tdr-021], we expect the information of up to 128 particles to be available in the trigger system [SCRecoL1Jet] after the phase-2 upgrade. However, the number of particles required for downstream tasks, varies depending on the application. For instance, while it has been shown that a good performance in b-tagging requires only $\sim 10$ particles [L1BTag], we have observed that the jet-tagging task on this dataset benefits from a larger number of particles in the input. Also, event-level tasks, such as missing transverse energy estimation or anomaly detection, may benefit from a larger input to represent the full event. To account for these variations, we train the models with 16, 32, 64, and 128 particles to cover the expected range of inputs.
@@ 3.2 Model
We adopt the MLP-Mixer architecture proposed in [mlp-mixer] as the backbone of our model. However, since the original model is too large for practical FPGA implementation, we designed a smaller version with only two mixer stages and a dense classification head. Nevertheless, we find that this simplified model is still capable to achieve state-of-art-performance. The exact architecture is depicted in Figure II. Following [jedi-fpga], we use all 16 particle features available in the dataset, and train the models with varying number of particles: 16, 32, 64, and 128.
@@
In contrast to [jedi-fpga], our models are not permutation-invariant, and all inputs are ordered by $p_{T}$. This design choice is made for several reasons:
@@
Figure II: Architecture of the MLP-Mixer models used in this work. Each model consists of four MLP blocks with a single skip-connection. The input channel size match the number of per-particle input features. The implementation of each MLP and the classification head is shown in the corresponding dashed blocks. MLP1 and MLP3 act on the feature dimension; MLP2 and MLP4 act on the particle dimension. DenseBn represents a dense layer followed by a batch normalization layer during training, which are fused into a single layer during inference. The exact implementation of the network can be found in the in the repository in Section Data availability statement.
@@ 3.3 Training
All models are implemented and trained with TensorFLow 2.13. We use the Adam optimizer with an initial learning rate of $5\cdot 10^{-3}$, along with the CosineDecayRestarts scheduler that resets the learning rate every 100 epochs. The full-precision models are trained for 500 epochs with a batch size of 512 with the categorical cross-entropy loss, and all models are trained on a single NVIDIA GTX 1080 GPU with 8GB of memory. We use the model with the best validation accuracy for performance evaluation described in Section 5.1 for the full-precision models.
@@ 4.1.1 High Granularity Quantization
We adopt the HGQ [hgq] method to train, quantize, and prune the MLP-Mixer models. HGQ is an advanced QAT method that optimizes the bitwidths of each individual parameter of the neural network with surrogate gradients on bitwidths. Comparing to traditional QAT methods that optimize the bitwidths at the layer level, such as QKeras [qkeras] or Brevitas [brevitas], HGQ provides a more fine-grained optimization that can achieve significant resource usage reduction for a fully unrolled firmware design while maintaining the accuracy of the original full-precision model. As HGQ allow the bitwidths to reach zero, unstructured pruning is naturally provided. HGQ is a crucial component of our workflow, which ensures that the MLP-Mixer models can be successfully deployed on FPGAs, as synthesis becomes infeasible if the model is not sufficiently quantized.
@@ 4.1.2 Distributed Arithmetic
DA [hcmvm, two-term, num-split, mmcm, fir1] is a method that transforms a linear digital signal processing (DSP) problem with one constant term, such as finite impulse response filters, discrete sine transform, or discrete cosine transform, into a series of operations implemented without the need for dedicated multipliers. This typically involves decomposing constant multiplication operations into a sequence of additions/subtractions and bit-shifts. Optimizations can be applied to the DA implementation to reduce the number of operations and the resource usage. In this work, we adopt a variant of the constant matrix-vector multiplication (CMVM) algorithm proposed in [two-term] to optimize the CMVM operations in the MLP-Mixer models. An outline of the optimization is as follows:
@@
We use the da4ml [da4ml-repo] library as an extension to the hls4ml library to apply this optimization. By applying the stated DA optimization, we significantly reduce the latency and initiation interval of the models.
@@ 4.2 Training
Following [hgq], we use a single training for each model to cover the entire Pareto frontier between accuracy and resource consumption [^1] The set of models that achieve the best accuracy for a given resource consumption, or the lowest resource consumption for a given accuracy. for the quantized models. We initialize the models with a small $\beta$ – a scalar parameter that encourages smaller models in a hardware resource-aware approach though regularization on differentiable bitwidths – and increase its value during training with an exponential scheduler. In each training step, the approximated Effective Bit-Operations ($\overline{\mbox{EBOPs}}$[hgq] – a differentiable resource surrogate when deployed on FPGAs for the model computed from the bitwidths from its parameters – is scaled by $\beta$ and added to the loss. In this way, $\beta$ leads to lower bitwidths, and consequently lower resource utilization on FPGAs. More details on $\beta$ and $\overline{\mbox{EBOPs}}$ can be found in [hgq]. During training, we save all models that lie on the Pareto frontier between validation accuracy and the resource consumption estimated by $\overline{\mbox{EBOPs}}$. All other hyperparameters remain identical to those used in the full-precision training. Each model is trained for 5,000 epochs to cover the Pareto frontier. However, the training time can be significantly reduced if only one performing model is required. For instance, all best-performing quantized models depicted in this work are obtained within 400 epochs.